OpenStudy (anonymous):

PLEASE HELP WILL MEDAL AND FAN Factor completely 5a2 + b. A. Prime B. a(5a + b) C. b(5a2) D. ab(5a + 1)

OpenStudy (anonymous):

Well you can't factor that, which means it's prime. For the other 3, multiply them out and you'll see that they don't equal 5a^2+b. a(5a + b) = 5a^2+ba b(5a2) = 5a^2b ab(5a + 1) = 5a^2b+ab

OpenStudy (anonymous):

thx @bobtheblob1416 <3